3.

Income Statement Sales Costs Balance Sheet Current 3,900 liabilities assets Current 2,100 assets 8,600 Long-term 3,700 5,500Fixed debt Taxable income $2,400 Equity 6,700 Taxes (25%) 600 Total $12,500 Total $12,500 Net income $ 1,800 Assets, costs, and current liabilities are proportional to sales. Long-term debt and equity are not. The company maintains a constant 40 percent dividend payout ratio. As with every other firm in its industry, next years sales are projected to increase by exactly 15 percent. What is the external financing needed? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

New sales = 7900*115% = 9085

New Costs = 5500* 115% = 6325

Taxable income = 9085 - 6325 = 2760

Tax@25% = 690

Net income = 2760 - 690 = 2070

Dividend payout is at 40%, plowback is at 60%

Plowback@60% = 2070 * 60% = 1242

Previous equity was 6,700, after plowback addition new equity will be 6700 + 1242 = 7942

New total assets = 12500 * 115% = 14375

New current liabilities = 2100 *115% = 2415

External financing = Total assets - total equity - total liabilities

=14375 - 7942 - 2415 - 3700

=$318